# PositionIC: Unified Position and Identity Consistency for Image Customization

### v2

@@ Abstract
Recent subject-driven image customization has achieved significant advancements in fidelity, yet fine-grained entity-level spatial control remains elusive, hindering the broader real-world application. This limitation is mainly attributed to the absence of scalable datasets that bind identity with precise positional cues. To this end, we introduce PositionIC, a unified framework that enforces position and identity consistency for multi-subject customization. We construct a scalable synthesis pipeline that employs a bidirectional generation paradigm to eliminate subject drift and maintain semantic coherence. On top of these data, we design a lightweight positional modulation layer that decouples spatial embeddings among subjects, enabling independent, accurate placement while preserving visual fidelity. Extensive experiments demonstrate that our approach can achieve precise spatial control while maintaining high consistency in image customization task. PositionIC paves the way for controllable, high-fidelity image customization in open-world, multi-entity scenarios and will be released to foster further research.
Figure 1: Rusults from PositionIC across various controllable image customization tasks.
@@ Introduction
Recently, diffusion-based models have achieved remarkable advances, especially in text-to-image (T2I) generation (Podell et al. 2023; Cai et al. 2025; Rombach et al. 2022; Li et al. 2024, ; Labs 2024). These models can produce images with superior visual quality and high semantic fidelity given short or descriptive prompts.
Among the tasks of image generation, image customization generation has attracted considerable attention in recent years. It requires models to generate images that not only adhere to textual descriptions but also preserve the identity features of reference objects. Although recent methods (Wu et al. 2025; Mou et al. 2025; Tan et al. 2024; Wang et al. 2025) have achieved notable improvements in image fidelity, they still lack customized control over the position, size, and scale of generated subjects and suffer from the limitation of insufficient high-quality customized paired data. While other methods (Wang et al. 2024c; Li et al. 2023) permit the placement of specific objects at specific positions, their position controllability and consistency are still insufficient and underexplored. This limitation hinders the effectiveness in applications where precise spatial arrangement and accurate visual quality are demanded, such as e-commerce product display, storybook illustration, and interior arrangement. As a result, there remains pressing challenges for approaches that can simultaneously ensure high image fidelity and offer flexible, fine-grained control over image attributes.
In this paper, we propose an automatic data curation pipeline and PositionIC, a novel method for controllable image customization to tackle these two challenges, i.e., (1) lack of high-quality customized paired data with position signals; and (2) controllable subject-driven customization guided by the layout. Specifically, regarding the data challenge, we introduce a Bidirectional Multi-dimensional Perception Data Synthesis (BMPDS) pipeline. We utilize a hierarchical training strategy to build a bidirectional generation paradigm from single-object to multi-object and vice versa, which could lift resolution constraints. We adopt Multi-modal Large Language Models (MLLMs) on data filtering to improve data quality. However, MLLMs are prone to visual hallucinations (Han et al. 2024; Pi et al. 2024), which can lead to inaccuracies when comparing visual information. Instead of directly relying on MLLMs to assess the consistency between two images, we carefully design a data filter that leverages expert models to convert visual information into textual descriptions. These descriptions are then used to assist MLLMs in performing multi-dimensional consistency evaluation. For controllable image customization, we propose PositionIC, a lightweight position-controllable framework that decouples layout from subject. It unlocks the fine-grained position control capability of Diffusion Transformers (DiT) through instance- level attention. Additionally, PositionIC does not introduce any additional parameters thus no additional computation is required. By restricting reference image features to focus only on specific regions, PositionIC can generate customized subjects at desired locations, which greatly expands the applications of controllable image customization as shown in Figure 1. Further experiments demonstrate the outstanding universality and remarkable capability of our method.
@@
In addition to using text prompts for conditional image generation, current diffusion models (Wei et al. 2023; Zhang et al. 2024; Huang et al. 2024; Xiao et al. 2025; Huang et al. 2024) support reference image input to achieve preservation of subject identity. Dreambooth (Ruiz et al. 2023) and LoRa (Hu et al. 2022) control the generation of diffusion models through fine-tuning on the specific subject. IP adapter (Ye et al. 2023) takes the reference images as feature inputs and achieves image-to-image generation by injecting the feature into the cross-attention module of the source model. Recently, DiT-based subject-driven models (Wu et al. 2025; Tan et al. 2024, 2025; Mou et al. 2025; Labs et al. 2025) further advance subject-driven generation. They introduce reference images as context information through token concatenation to ensure the consistency of objects during generation process.
@@
Position-controllable image-driven customization requires high-fidelity paired data, featuring prominent subjects and high resolution with layout control signals. However, existing open-source datasets such as Subject200K generate paired data using diptych images, which leads to object inconsistency issues and are limited by low resolution and the lack of positional information. We introduce BMPDS, a Bidirectional Multi-dimensional Perception Data Synthesis framework to tackle these limitations. We adopt a hierarchical generation-and-selection strategy to progressively improve data quality, gradually introducing single-image and multi-image pairs with spatial control information during generation. The overall framework is depicted in Figure 2.
@@
We divide the automated data generation process into three stages. (1) Inspired by UNO, we first train a weak model for image customization tasks using Subject200K (Tan et al. 2024) dataset. The generated results are then segmented and passed through a Flux-Outpainting model with random placement to inject spatial control information. The paired data obtained in this stage is filtered by a filter to ensure high fidelity. These data are then used to train the PositionIC-Single model. (2) In the second stage, we perform forward generation of multi-image data pairs. We independently process Subject200K samples via PositionIC-Single, then randomly pair and position the output as input to the Flux-Outpainting model, thereby obtaining multi-subject paired data. (3) To enhance data diversity and improve generalization, we reverse the above process in the third stage. We first use Multi-modal Large Language Models (MLLMs) to generate text descriptions containing multiple subjects, then employ the Flux model to create high-resolution images. Objects are detected and cropped from these results, individually processed by the PositionIC-Single, resulting in high-resolution multi-subject paired data.
Through bidirectional data synthesis, we construct PIC-400K dataset with noise interference.
@@
Previous studies have shown that MLLMs have limited capability in recognizing fine-grained details in images. Instead of directly feeding image pairs into the MLLMs for filtering, we establish a reliable system to achieve more accurate and efficient data filtering after data generation. Specifically, we divide the filtering process into three levels based on granularity. Firstly, we segment the subjects from data pairs and utilize CLIP-I (Radford et al. 2021) and DINO (Caron et al. 2021) scores $s_{v}$ to filter out images with significantly lower consistency. Subsequently, we pass two subjects’ images through MLLMs (e.g., GPT-4o), which directly gives a similarity score $s_{vlm}$ based on shape, color, and details. Lastly, we employ Describe Anything Model (DAM) (Lian et al. 2025), a description expert to obtain detailed textual descriptions for each subject. Given these textual description pairs, we instruct GPT-4o as a judge to autonomously select comparative features(e.g., color, shape) and assign multi-dimensional similarity scores $s_{ds}$.
@@
We apply the filter on PIC-400K to rank the pairs and filter out inconsistent pairs. The filtered dataset PIC-98k consists of 44k single-subject pairs and 54k multi-subject pairs.
@@
To manage subject consistency generation position simultaneously (e.g., generating the subject in a specific location in the image), we introduce a light-weight approach named attention accumulating, which can unlock DiT’s ability of spatial control without extra training and inference cost. As shown in Figure 3, the attention map can be divided into four areas: text-text self attention, image-image self attention, text-image and image-text attention. If extending to single or multi-subject custom generation via concatenating method like (Wu et al. 2025; Tan et al. 2024; Labs et al. 2025; Zhang et al. 2025), the attention map will expand to nearly three times its original size, making it harder for the model to focus on the corresponding regions.
To achieve effective attention accumulating, we explicitly define the area that can be focused on for each reference image, which we name Attention Horizon. Previous works (Zhang et al. 2025; Chen et al. 2024) explore the effort of restricting the attention horizon between special words and noise, we extend it to subject-driven generation task to unlock the positional control ability of diffusion transformers (DiT).
@@
To evaluate our model’s performance on the benchmark, We leverage multi-modal large language models (MLLMs) to detect objects and assess their content, computing the IoU and object category alignment between the generated and reference subjects.
Figure 7: Qualitative comparison of position control generation with different methods on PositionIC-Bench. RPF denotes Regional Prompting Flux. MS-Diff and Gligen are existing position-controllable customization methods; RPF and Instance-Diff are position-only controllable methods.
@@
DreamO 0.835 0.258 0.802
PositionIC$({Ours})$ 0.846 0.269 0.823
Table 1: Evaluation on DreamBench for single-subject driven generation. The bold value is the best and the underlined value is the second best.
@@
UNO 0.781 0.279 0.707
DreamO 0.779 0.273 0.698
PositionIC$({Ours})$ 0.819 0.279 0.771
@@
For the position-guided task, we evaluate different methods on our proposed PositionIC-Bench We use Vision-R1 (Huang et al. 2025) to determine the bounding box of the subject and calculate mIoU and AP scores with the label.
@@
MS-Diffusion 0.501 0.097 / 0.329 / 0.075 0.421 0.028 / 0.146 / 0.005
Instance-Diffusion 0.789 0.593 / 0.683 / 0.632 0.799 0.497 / 0.699 / 0.546
PositionIC$({Ours})$ 0.828 0.628 / 0.904 / 0.761 0.860 0.701 / 0.939 / 0.853
@@
We visualize the comparison results with the current state-of-the-art methods at Figure 4. Overall, our method surpasses all current methods in terms of visual effects. It can be seen that our method can still effectively follow the prompt while maintaining subject consistency, demonstrating higher text fidelity. Other methods either fail to follow complex instructions or cannot maintain consistency. (e.g., UNO and DreamO cannot reproduce the dog face, while SSR-Encoder fails to add the Santa hat.) The results of the third row also reveal that PositionIC has a great capability on patterns and text. Furthermore, PositionIC consistently produces images with a higher degree of naturalness and visual plausibility.
Figure 5 shows the comparison of multi-subject generation. To control variables, the images generated by our method adopt a fixed bounding box (e.g., bottom left and bottom right). It can be seen that in the more difficult multi-subject scenario, PositionIC can still maintain high subject similarity and follow the given text prompt, whereas results from most other methods fail to preserve consistency for each subject and even ignore certain subjects.
@@
Specifically, we discover that different object sizes can lead to inaccurate scores. As shown in Figure 6, . To avoid the sensitivity, we crop the subject from original image as source images for evaluation.
@@
In this section we show the ablation study of our key modules, including horizon mask and data quality. The results are shown in Figure 8 and Table 4.
Impact of data filter. Results in Figure 8 illustrate the advanced fidelity of BMPDS. We directly trained with PIC-98K, PIC-400K and the Subject200K dataset without injecting position control information. PIC-400K achieves significantly higher CLIP-I scores than Subject200K, and the filtered data PIC-98K achieves the highest scores overall, which demonstrates the remarkable efficiency of our data synthesis and filtering pipeline.
@@
Horizon mask. As shown in Table 4, CLIP-I and DINO scores significantly drop when training without horizon mask. We believe that by adopting horizon mask, model can focus the transfer and generation of image features on a smaller region rather than globally, which accelerates the convergence and improve the consistentcy.

### v3

@@ Abstract
Recent subject-driven image customization has achieved significant advancements in fidelity, yet fine-grained instance-level spatial control remains elusive, hindering the broader real-world application. This limitation is mainly attributed to the absence of scalable datasets that bind identity with precise positional cues. To this end, we introduce PositionIC, a unified framework that enforces position and identity consistency for multi-subject customization. We construct a scalable synthesis pipeline that employs a bidirectional generation paradigm to eliminate subject drift and maintain semantic coherence. On top of these data, we design a lightweight positional modulation operation that decouples spatial embeddings among subjects, enabling independent, accurate placement while preserving visual fidelity. Extensive experiments demonstrate that our approach can achieve precise spatial control while maintaining high consistency in image customization task. PositionIC paves the way for controllable, high-fidelity image customization in open-world, multi-entity scenarios and will be released to foster further research.
@@ Introduction
Diffusion-based models have recently revolutionized visual synthesis, especially in text-to-image (T2I) generation, where they produce photorealistic images that faithfully match short textual prompts (Podell et al. 2023; Esser et al. 2024; Rombach et al. 2022; Li et al. 2024; Labs 2024).
Within this broader area, subject-driven image customization seeks to generate new scenes that simultaneously (i) conform to user-provided textual descriptions and (ii) preserve the identity of one or more reference objects. Although recent methods (Wu et al. 2025; Mou et al. 2025; Tan et al. 2024; Wang et al. 2025) markedly improve visual fidelity, they still offer limited control over where or how each subject appears (i.e., position, scale, and relative layout). This limitation stems from twofold root causes: scarce data with explicit positional labels and existing global-level attention mechanisms that cannot deliver fine-grained, instance-level guidance.
Figure 1: Results from PositionIC across various controllable image customization tasks.
Approaches that do permit object placement (Wang et al. 2024c; Li et al. 2023) often compromise either spatial accuracy or identity consistency, restricting their usefulness in practical scenarios such as e-commerce product display, storybook illustration, and interior design.
Consequently, it remains an open challenge to achieve both high-fidelity rendering and fine-grained, flexible control over subject attributes in a unified framework.
To this end, we address these issues with two tightly coupled innovations: i) Bidirectional Multi-dimensional Perception Data Synthesis (BMPDS), a scalable pipeline for generating high-quality, position-annotated multi-subject data; ii) PositionIC, a lightweight, layout-aware framework that enables fine-grained, position-controllable, identity-consistent image customization.
Specifically, to overcome the data bottleneck, we devise an automatic pipeline that expands single-subject collections into scalable, high-quality multi-subject datasets annotated with explicit position masks. A hierarchical training schedule establishes a bidirectional generation paradigm, progressively moving from single-to-multi subject synthesis and back, so that resolution constraints are relaxed while subject drift is suppressed. Because of the inherent hallucination of Multi-modal Large Language Models (MLLMs) (Han et al. 2024; Pi et al. 2024), we avoid direct visual comparisons: expert vision models first translate visual content into textual descriptions, after which MLLMs perform multi-dimensional consistency checks. This two-stage filtering markedly improves data reliability.
Built upon the BMPDS corpus, PositionIC injects an instance-level attention modulation into Diffusion Transformer (DiT), endowing its fine-grained position control capability. By decoupling instance-level spatial embeddings from semantic identity features, our method enables independent, accurate placement of each subject while introducing no extra train-time parameters or inference overhead. Restricting reference features to user-specified regions further enhances identity fidelity and spatial precision, unlocking new applications of controllable image customization (Figure 1).
@@
In addition to using text prompts for conditional image generation, current diffusion models (Wei et al. 2023; Li, Li, and Hoi 2023; Huang et al. 2024; Ye et al. 2023; Xiong et al. 2025; Xiao et al. 2023; Bar-Tal et al. 2023; Feng et al. 2025) support reference image input to achieve preservation of subject identity. Dreambooth (Ruiz et al. 2023) and LoRa (Hu et al. 2022) control the generation of diffusion models through fine-tuning on the specific subject. Recently, Diffisuion-Transformers-based subject-driven models (Wu et al. 2025; Tan et al. 2024, 2025; Mou et al. 2025; Xiao et al. 2025; Labs et al. 2025) further advance subject-driven generation. They introduce reference images as context information through token concatenation to ensure the consistency of objects during generation process.
@@
Position-controllable image-driven customization requires high-fidelity paired data, featuring prominent subjects and high resolution with layout control signals. However, existing open-source datasets such as Subject200K (Tan et al. 2024) generate paired data using diptych images, which leads to object inconsistency issues and are limited by low resolution and the lack of positional information. We introduce BMPDS, a Bidirectional Multi-dimensional Perception Data Synthesis framework to tackle these limitations. We adopt a hierarchical generation-and-selection strategy to progressively improve data quality, gradually introducing single-image and multi-image pairs with spatial control information during generation. The overall framework is depicted in Figure 2.
@@
We divide the automated data generation process into three stages. (1) Inspired by UNO, we first train a weak model for image customization tasks using Subject200K dataset. The generated results are then segmented and passed through a Flux-Outpainting model with random placement to inject spatial control information. The paired data obtained in this stage is filtered by a filter to ensure high fidelity. These data are then used to train the PositionIC-Single model. (2) In the second stage, we perform forward generation of multi-image data pairs. We independently process Subject200K samples via PositionIC-Single, then randomly pair and position the output as input to the Flux-Outpainting model, thereby obtaining multi-subject paired data. (3) To enhance data diversity and improve generalization, we reverse the above process in the third stage. We first use Large Language Models (LLMs) to generate text descriptions containing multiple subjects, then employ the Flux model to create high-resolution images. Objects are detected and cropped from these results, individually processed by the PositionIC-Single, resulting in high-resolution multi-subject paired data.
Through bidirectional data synthesis, we construct PIC-400K dataset. However, we need extra filtering processes to improve data quality as the dataset is still suffering from noise.
@@
Previous studies (Cao et al. 2024; He et al. 2025) have shown that MLLMs have limited capability in recognizing fine-grained details in images. Instead of directly feeding image pairs into the MLLMs for filtering, we establish a reliable system to achieve more accurate and efficient data filtering after data generation. Specifically, we divide the filtering process into three levels based on granularity. Firstly, we segment the subjects from data pairs and utilize CLIP-I (Radford et al. 2021) and DINO (Caron et al. 2021) scores $s_{v}$ to filter out images with significantly lower consistency. Subsequently, we pass two subjects’ images through MLLMs (e.g., GPT-4o), which directly gives a similarity score $s_{vlm}$ based on shape, color, and details. Lastly, we employ Describe Anything Model (DAM) (Lian et al. 2025), a description expert to obtain detailed textual descriptions for each subject. Given these textual description pairs, we instruct GPT-4o as a judge to autonomously select comparative features(e.g., color, shape) and assign multi-dimensional similarity scores $s_{ds}$.
@@
We apply the filter on PIC-400K to rank the pairs and filter out inconsistent pairs. The filtered dataset PIC-98k consists of 44k single-subject pairs and 54k multi-subject pairs. More examples of our PIC-98K are depicted in Appendices.
@@
To manage subject consistency generation position simultaneously (e.g., generating the subject in a specific location in the image), we introduce a light-weight approach named attention accumulating, which can unlock DiT’s ability of spatial control without extra training and inference cost. As shown in Figure 3, the attention map can be divided into four areas: text-text self attention, image-image self attention, text-image and image-text attention. If extending to single or multi-subject custom generation via concatenating method like (Wu et al. 2025; Tan et al. 2024; Labs et al. 2025; Zhang et al. 2025a), the attention map will expand to nearly three times its original size, making it harder for the model to focus on the corresponding regions.
To achieve effective attention accumulation, we explicitly define the area that can be focused on for each reference image, which we name Attention Horizon. Previous works (Zhang et al. 2025a; Chen et al. 2024; Zhang et al. 2025b) explore the effort of restricting the attention horizon between special words and noise, we extend it to subject-driven generation task to unlock the positional control ability of diffusion transformers (DiT).
@@
Figure 7: Qualitative comparison of position control generation with different methods on PositionIC-Bench. MS-Diff and RPF denote MS-Diffusion and Regional Prompting Flux respectively. MS-Diff and Gligen are existing position-controllable customization methods; RPF and Instance-Diff are position-only controllable methods.
@@
PositionIC(Ours)({Ours}) 0.846 0.269 0.823
Table 1: Evaluation on DreamBench for single-subject driven generation. The bold value is the highest and the underlined value is the second.
@@
PositionIC(Ours)({Ours}) 0.819 0.279 0.771
@@
For the position-guided task, we evaluate different methods on our proposed PositionIC-Bench. We use Vision-R1 (Huang et al. 2025) to determine the bounding box of the subject and calculate mIoU and AP scores with the label.
@@
PositionIC(Ours)({Ours}) 0.828 0.628 / 0.904 / 0.761 0.860 0.701 / 0.939 / 0.853
@@
We visualize the comparison results with the current state-of-the-art methods in Figure 4. Overall, our method surpasses all current methods in terms of visual effects. It can be seen that our method can still effectively follow the prompt while maintaining subject consistency, demonstrating higher text fidelity. Other methods either fail to follow complex instructions or cannot maintain consistency. (e.g., UNO and DreamO cannot reproduce the dog face, while SSR-Encoder fails to add the Santa hat.) The results of the third row also reveal that PositionIC has a great capability on patterns and text. Furthermore, PositionIC consistently produces images with a higher degree of naturalness and visual plausibility.
Figure 5 shows the comparison of multi-subject generation. To control variables, the images generated by our method adopt a fixed bounding box (e.g., bottom left and bottom right). In a more difficult multi-subject scenario, PositionIC can still maintain high subject similarity and follow the given text prompt, whereas results from most other methods fail to preserve consistency for each subject and even ignore certain subjects.
@@
Figure 8: User study for subject-driven generation and data filter. (a) User evaluation on Dreambench. (b) Filtering consistency of our data filter compared with human in different data quality intervals (from good to bad).
Quantitative Evaluations
@@
Specifically, we discover that different object sizes can lead to inaccurate scores. As shown in Figure 6, to avoid the sensitivity, we crop the subject from original image as source images for evaluation.
@@
User Study
We invite evaluators for an extensive user study. For subject-driven generation, we randomly selected 500 images from the results on DreamBench for manual evaluation. There are six users scored the results from five dimensions, with each ranging from 0 to 5, and the average score was taken. The result presented in Figure 8 (a) shows that our method reaches the highest capability to preserve image features and details while maintaining competitive text adherence.
To evaluate the consistency between human annotators and the data filter in BMPDS, we use percentage agreement metric (denoted as Acc.), comparing the filter’s output against human-generated annotations. As shown in Figure 8 (b), our filter has an average consistency of 0.89 with human annotators.
@@
In this section we show the ablation study of our key modules, including horizon mask and data quality. The results are shown in Figure 9, Table 4 and Figure 10.
Impact of data filter.
Results in Figure 9 illustrate the advanced fidelity of BMPDS. We directly trained with PIC-98K, PIC-400K and the Subject200K dataset without injecting position control information. PIC-400K achieves significantly higher scores than Subject200K, and the filtered data PIC-98K achieves the highest scores overall, which demonstrates the remarkable efficiency of our data synthesis and filtering pipeline.
@@
Figure 10: Qualitative results of ablation study. Horizon mask and our filter pipeline are capable of effective position control and consistent customization.
Horizon mask.
As shown in Table 4, CLIP-I and DINO scores significantly drop when training without horizon mask. We believe that by adopting horizon mask, model can focus the transfer and generation of image features on a smaller region rather than globally, which accelerates the convergence and improve the consistency.
@@ Conclusion
Appendix A Bidirectional Multi-dimensional Perception Data Synthesis
Detailed Instruction of GPT-4o
As shown in Figure 14, the message given to GPT-4o consists of Instruction, Evaluation Metric and Response. In the part of Instruction, we have defined the input format and evaluation metric dimensions, and required GPT-4o to select no fewer than three features for scoring based on the content of textual description. In the next part of Evaluation Metric, we detail the metric standard and provide GPT-4o examples to evaluate. There are 6 levels, ranging from 0 to 5, representing the similarity between two descriptions regarding the same feature. After that, we prompt GPT-4o to return a dictionary in JSON format containing the subject types and the scores for each feature. If there is no similarity between the two descriptions, the subject is set to ”none”, indicating that the final score is 0.
Due to the substantial differences in textual descriptions between subjects, it is not feasible to predetermine the feature categories for evaluation. Therefore, we allow the LLMs to select at least three features and assign individual scores to each. The final score is calculated as the average of all feature scores. For descriptions with significant discrepancies, the LLMs is permitted to assign a score of zero to the samples.
Figure 15 demonstrates the samples of Multi-dimensional Perception Data Filter. We have highlighted the correlated features in the description. In the first sample, the teddy bear share the same physical characteristics except for their posture, hence earning the highest appearance score and slightly lower posture score. In the second sample, the deer is missing antlers, which resulted in the lowest score on the ”antler” feature.
Details of PIC-98K Dataset
We propose PIC-400K utilizing our Bidirectional Multi-dimensional Perception Data Synthesis, a automatic and effective high-consistency data synthesis pipeline. Samples of the filtered data PIC-98K is shown in the Figure 13. BMPDS can synthesize high-fidelity multi-subject images while maintaining high resolution. Against previous works, the position of subjects is controllable and it is randomly placed to train the position control capability of PositionIC.
There are over 9000 subject descriptions in PIC-98K, including multiple categories such as fruits, animals, and transportation vehicles, which basically cover common objects. The distribution of subjects is shown in Figure 11, vehicles, furniture, animal, and kitchenware constitute a significant proportion, with most difficult-to-classify subjects categorized as ”other”.
Figure 12: Showcases of PositionIC-Bench.
Appendix B PositionIC-Bench
We manually select 252 single-subject samples and 296 multi-subject samples in the benchmark, where the object bounding boxes conform to standard proportions and include challenging positional relationships. We show some samples of PositionIC-Bench in Figure 12. Our bench includes various subjects such as furniture, animals, plants, and portraits. Not limited to conventional object placement, PositionIC-Bench’s bounding boxes have more complex spatial relationships where objects are placed on different planes. At the same time, the bounding boxes of smaller objects is appropriately enlarged to obtain more accurate evaluation scores.
Figure 13: Showcases of PIC-98K Dataset.
Figure 14: Prompt template of MLLMs in Multi-dimensional Perception Data Filter.
Figure 15: Examples of Multi-dimensional Perception Data Filter.